OpenStudy (anonymous):

Find the derivative of the function using the definition of derivative. f(x) = (1/2)x-(1/6) I found that f'(x) =1/2 I don't know what the domain of the function, f(x) is and what the domain of its derivative, f'(x) is.

OpenStudy (anonymous):

Domain of f(x)= all real nos.

OpenStudy (anonymous):

\[f \left( x+h \right)=\frac{ 1 }{2 }\left( x+h \right)-\frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

find f(x+h)-f(x) and divide by h,and finally take limit as h approaches 0

OpenStudy (anonymous):

I solved but my connection failed.